OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (saifoo.khan):

Options?

OpenStudy (anonymous):

y + 3 = 2(x – 4) y – 3 = 2(x + 4) y + 4 = 2(x – 3) y – 4 = 2(x + 3)